answersLogoWhite

0

🎒

Statistics

Statistics deals with collecting, organizing, and interpreting numerical data. An important aspect of statistics is the analysis of population characteristics inferred from sampling.

36,755 Questions

What dose sequence of events mean?

The logic progress of certain events and how they are linked together.

To cut the maximum likely error in half the sample size should be?

... should be increased by a factor of 4. Note that this implies that the only errors are statistical (random) in nature; increasing the sample size won't improve systematic errors.

Anuzinc suppository dose frequency?

"Directions: Remove wrapper and insert one suppository into the rectum in the morning and at bedtime, and after each bowel movement."

A negative correlation was found between weight and distance walked. Would you conclude that walking causes lower weight and can you think of an alternate explanation for the findings?

Negative correlation implies an inverse relationship between a person's weight and the amount of distance that they walk (say on an average day). So you can conclude that heavier people are more likely to walk less. Lighter people are more likely to walk more.

Remember though that correlation does not always equate to causation. What this means in practical terms is that walking more does not necessarily mean you'll lose weight. Conversely, walking less does not necessarily mean that you'll gain weight. Or thinking of it another way, losing weight will not necessarily cause you to walk more and gaining weight will not necessarily cause you to walk less.

How many numbers does the creditcard have?

16, in four sets of four. Plus expiry dates, and security numbers at the back.

2 tons is equal to how much pounds?

There are 2000 pounds in a ton; therefore 2 tons is 4000 pounds.

If a diameter of the semicircle is 14cmfind its perimeter?

Perimeter = Circumference/2+Diameter Perimeter = (pi*14)/2+14 = 35.99114858cm or 36cm correct to two significant figures.

How do you find the medium mean and mode in a set of numbers?

First arrange the numbers in ascending order

The middle number is the median

Add the numbers together and divide them by how many number there are for the mean

The number that occurs most often is the mode

What is the probability of getting a license plate that has a repeated letter or digit if you live in a state that has one numeral followed by six letters followed by one numeral?

Probability of getting a repeated digit = no. of favourable outcomes/total no. of possible outcomes

Favourable outcomes=(0,0),(1,1),(2,2).....(9,9)

thus no. of favourable outcomes = 10

Considering that anyone of the 10 digits may apperar as the first numeral as well as d last numeral,

No. of possible outcomes=10*10=100

hence probablity of a repeated digit=10/100=0.1

Probability of getting a repeated letter = no. of favourable outcomes/total no. of possible outcomes

consider 6 blanks _ _ _ _ _ _ ,each filled with a letter.

Thus no. of possible outcomes = 26^6

now consider that any two of these blanks have the same letter.

Consider the two blanks filled with same letter as one blank.

_ _ _ _ _

So that blank can be filled in 26 ways(i.e. you can have any of 26 alphabets as the repeated letter)

The other 4 blanks can be filled with rest of 25 alphabets as the one that has already been used(the repeated letter) cannot be used again. We want all other letters to be differrent. So the next four blanks can be filled in 25*24*23*22ways.

Thus,

no. of favaourable outcomes = 26*25*24*23*22

Probability of getting a repeated letter=(26*25*24*23*22)/(26^6)

=0.0255(approx)

And,

Total probablity of getting A repeated digit OR letter=0.1+0.0255=0.1255

i.e. 12.55%

__________

Another approach, if the letters are able to be recurring throughout, which they usually are, you have a 1 in 95,428,956,661,682,176 of getting two of the same letters right next to each other.

You first take 26^6, then it comes out as a 1 in 308,915,776 of getting a letter you want. Then you square that, because the outcome MUST be the same, so the chances of you getting that same letter increase by whatever the denominator is, which would be 308,915,776^2. Hopefully I'm right. xD

Which graph compares data from a number of distinct categories?

In general, graphs compare two variables. A scatter plot can compare a third variable because the quantity of dots gives additional information from just the two axes. A color coded scatter plot can give even more information because each colored dot can be distinct.

Adding more colors can ad more distinct categories.

Color coding any type of graph, such as pie chart, bar graph or multi-line two axis graphs can also represent more than two distinct categories.

Why with random number repeat previous number often?

Given any random digit, there is a 10% probability that the next digit is the same. That is, the probability of the previous number repeating is 10%. Whether this constitutes often (as suggested by the question) or not is a matter of opinion.

How many digits are there in all numbers inclusive from 1 to 1000?

Infinitely many. The number pi , for example, is between 1 and 1000 and, since pi is a transcendental number, it contains infinitely many digits. Plus, there are all the irrational numbers - each with infinitely many digits, and all the rationals with recurring decimals - again with infinitely many digits.